# Iñigo Alonso1, Imanol Miranda2, Eneko Agirre2, Mirella Lapata1

### v2

@@ 3.4 Linking to Sources and Highlighting
Tablet is designed with extensibility in mind. Existing datasets such as TableInstruct (zhang-etal-2024-tablellama), DocStruct4M (hu-etal-2024-mplug), and MMTab (zheng-etal-2024-multimodal) do not provide a clear reference to the original examples from which their instructions were derived. However, the absence of such pointers, makes reuse, modification, and augmentation difficult. Additionally, image-based datasets such as MMTab lack serialized table versions in text format, providing only rendered images as representations. To address these issues, each example in Tablet includes both the identifier of the original example and the identifier of the corresponding table from the source dataset. Identifiers follow a simple, human-readable format, and the released code provides a function to retrieve the original examples given their IDs.
To support future extensions and new tasks, Tablet also provides serialized versions of its tables in HTML. For the tables successfully retrieved from Wikipedia, this HTML corresponds to the raw table object in the original HTML of the article. These tables can be rendered with their full visualization using Wikipedia’s official CSS stylesheets or through the rendering functions provided with our code. The remaining tables are either provided as HTML in the seed datasets or converted into HTML from their serialized representations.
This feature is particularly useful in tasks involving cell pointing or highlighting. Prior work has shown that models can achieve competitive results by relying uniquely on highlighted values (an2022cont; alonso-etal-2024-pixt3). We also observed in our experiments that as long as highlighted cell values are mentioned in the prompt, models are able to perform the task (e.g., Column Type Annotation, Entity Linking, Relation Extraction, FeTaQA, and ToTTo) regardless of the table provided. In other words, the model learns to ignore the table during training, which can be detrimental for tasks requiring genuine table understanding. We mitigate this, by exploiting the HTML serialization to directly locate the highlighted cells in the source tables. By cross-referencing the highlighted spans from the seed datasets with the actual table content, we generate versions of the tables with explicit highlights while preserving the original visual features.
@@ 3.5 Comparison with MMTab
MMTab (zheng-etal-2024-multimodal) is also a dataset for visual table understanding; however, unlike Tablet, it does not include the full set of training examples from the original datasets (e.g., only 12.4% of ToTTo compared to our 91.8%; see Appendix 7.6), and it provides no development splits. MMTab covers 20 tasks while Tablet covers 21 (inculding instruction tuning); four tasks are unique to the former and five to the latter, and overall our dataset is 9.36 times larger (comprising 4,066,545 examples compared to MMTab’s approximate 433,376 examples). Moreover, only a few tasks in MMTab include identifiers that can be linked back to their source examples, limiting developers to the provided instruction text. While MMTab relies on synthetically rendered tables with predefined styles, which omit original formatting and embedded images, Tablet retains the original visualizations, enabling models to exploit a fuller range of visual features.
@@ 4 Experimental Setting
Our experiments are motivated by two questions: (1) Does training VTU models with lossless table visualizations improve performance on VTU tasks? and (2) Is Tablet diverse enough so that supervised fine-tuning (SFT) improves generalization on unseen VTU tasks? We also examine whether training only with original visualizations performs better than mixing original and synthetic ones, whether a balanced task distribution is preferable to the full dataset distribution, and whether including various table interpretation tasks contributes to VTU performance.
Backbone Model All experiments employ the same backbone model, namely Qwen2.5-VL with 7B parameters (qwen2025qwen25technicalreport). This model provides a good tradeoff between performance and computational requirements. Together with InternVL3 (zhu2025internvl3exploringadvancedtraining), it was the best-performing open-weight VLM in its class, setting a new benchmark for Single Page Document VQA (mathew2021docvqa) at the time of writing. Exploring the best backbone model for VTU or advanced SFT strategies is outside the scope of this work. Full SFT configurations are detailed in Appendix 8; results for other open-weight models on Tablet (test set) are reported in Appendix 8.4.
Evaluation We evaluate fine-tuned models on eight held-in tasks: three focus on Table QA (WikiTableQuestions, HiTabQA, FeTaQA), two on text generation (WikiBio, ToTTo), two on Table Numerical Reasoning tasks (TabMWP, TAT-QA), and one on Table NLI (TabFact). To test generalization, we further evaluate on six held-out tasks: HybridQA, InfoTabs, TabMCQ (jauhar2016tabmcqdatasetgeneralknowledge), AIT-QA (katsis-etal-2022-ait), Table Recognition, and PubHealthTab (akhtar-etal-2022-pubhealthtab). The last four are directly extracted from MMTab and use instructions generated with their own templates, adding diversity in both table and instruction styles. Although included in Tablet, the 86,135 training examples from HybridQA, InfoTabs, and Table Recognition were excluded from fine-tuning to enable held-out evaluation. Models trained on the full dataset would potentially achieve even stronger results than those reported here.
5 Results and Analysis
[…]
WikiBio ToTTo WikiTQ TabMWP HiTAB TabFact FeTaQA TAT-QA
Model synth org synth org synth org org synth org synth org synth org org
0-Shot 2.0 2.5 10.8 8.9 57.7 50.8 51.9 11.1 2.5 70.4 59.4 9.9 9.8 48.7
@@ 5 Results and Analysis
Table 2: Comparison of Qwen2.5-VL-7B-Instruct models fine-tuned on different Tablet-B variants and in zero-shot mode; models are evaluated on original (lossless) table visualizations and synthetic ones across held-in benchmarks. Results for TabMWP and TAT-QA exclude the mixed setting, as all tables already use original visualizations. We report exact match accuracy for WikiTQ, TabMWP, TabFact, and TAT-QA; BLEU for WikiBio, ToTTo, and FeTaQA; and F1 for HiTab. ToTTo results are reported on its development set. Best model per setting (synth/org) shown in bold.
Our results in Table 2 (held-in tasks) show that lossless models outperform synthetic ones in two tasks (WikiBio, HiTab), match performance in four (WikiTQ, TabMWP, FeTaQA, TAT-QA), and underperform in one (TabFact). For reference, we also report the 0-shot performance of Qwen2.5-VL-7B. The fact that there is no pronounced difference between lossless and synthetic visualizations is perhaps not surprising since most benchmarks were originally designed around serialized, text-only representations, without paying heed to visual cues (TURL, WikiBIO, ToTTo, TAT-QA). Thus, while lossless visualizations contain more information, current benchmarks do not test whether models benefit from them. Developing benchmarks with high visual variability and challenging layouts would better highlight scenarios where lossless visualizations offer a distinct advantage.
@@
Performance on Unseen VTU Tasks with Tablet We next evaluate whether Tablet’s task diversity improves performance on unseen datasets. While fine-tuned models are expected to perform better on tasks seen during training, improved performance on held-out datasets would suggest that training on Tablet enhances the model’s VTU capabilities. Tables 3 and 4 report our results on held-in and held-out tasks, respectively. We present comparisons between zero- and one-shot Qwen2.5-VL-7B [^4] We should note that Qwen2.5-VL-7B is a competitive baseline that has undergone extensive pre-training and instruction-tuning, achieving top scores in many TU and Visual Document Understanding benchmarks. and its fine-tuned instantiations trained on MMTab and different sizes of Tablet. For MMTab, all Wikipedia-sourced images are synthetic (61.8% of all tables in their dataset), whereas all Tablet versions use a mix of original (88%) and synthetic (12%) (See Appendix 7.2 for the distribution of image sources per task). For now, we focus solely on Tablet-L (a shorthand for Tablet-Large), which is our biggest dataset comprising 4M examples (we discuss smaller sizes in the next sections).
@@
Table 3: Evaluation on eight held-in datasets  with Qwen2.5-VL-7B-Instruct in 0/1-shot mode and fine-tuned on MMTab and different Tablet sizes. Results for ToTTo correspond to evaluation on the development set. We report exact match accuracy for WikiTQ, TabMWP, TabFact, and TAT-QA; BLEU for WikiBio, ToTTo, and FeTaQA; and F1 for HiTab. Best performing model per task is shown in bold; we mark with $\ast$ models significantly different ($p<0.05$, using bootstrap resampling) from those trained on Tablet (highlighted in gray). Models that perform significantly better within the Tablet group are underlined.
@@
Infotabs TabMCQ AIT-QA PubHealthTab HybridQA TabRec
0-Shot  64.5 84.4  51.7  64.7 34.2  24.5
1-Shot 65.5  61.0  43.1  70.0  30.5  27.9
MMTab  56.9 89.1  56.6  63.9 27.1  43.6
@@
Table 4: Left: Evaluation on five held-out datasets with Qwen2.5-VL-7B-Instruct (0/1-shot) and fine-tuned on MMTab and different Tablet sizes. We report Tree-Edit- Distance-based Similarity for Table Recognition (TabRec) and accuracy for all other benchmarks. HybridQA results are reported on the development set. Best performing model per task is shown in bold; we mark with $\ast$ models significantly different ($p<0.05$, using bootstrap resampling) from those trained on Tablet (highlighted in gray). Models that perform significantly better within the Tablet group are underlined. Right: Distribution of examples per task in Tablet-L (b) and Tablet-M (a).
@@
We observe that the model fine-tuned on Tablet-L  performs overwhelmingly better on held-in and held-out datasets. It dominates in all QA benchmarks except HybridQA, where long-context proficiency is likely reduced since its training set was excluded from fine-tuning. Notably, in tasks such as HiTab, FeTaQA, and TAT-QA, where Tablet includes fewer training examples than MMTab, the model still performs considerably better, suggesting benefits from transfer across tasks, original visualizations, or simply due to a higher-quality subset of examples. Finally, tasks that include hierarchical table structures such as ToTTo and HiTab show clear gains. Table cell hierarchy is not lost in synthetic tables, but it is often conveyed more clearly in their original visualizations. Our results are encouraging, particularly when considering that portions of the held-out datasets, including test sets, may have been encountered during petraining. Models trained on Tablet outperform 0/1-shot Qwen2.5-VL-7B on 12 out 14 (held-in and held-out) tasks.
Interestingly, our gains extend beyond unseen tasks. For instance, ToTTo, HiTab, TabFact, FeTaQA, and TAT-QA contain the same number of training examples across all Tablet variants, yet performance consistently improves as additional tasks are incorporated. This suggests that fine-tuning on a broader task set facilitates transfer learning, yielding improvements even on tasks where the model was already competitive.
@@
Optimal Tablet Size Since training on 4M examples is resource-intensive, and tasks in Tablet-L exhibit substantial variation in size (see (b) pie chart in Table 4), we evaluate whether a smaller but more balanced dataset might perform comparably. We create Tablet-M (a shorthand for Tablet-Medium) by capping each task at 140k examples. For benchmarks exceeding this cap (Column Type Annotation, Entity Linking, WikiBio), a random 140k subset is sampled and included as representative for that task. This results in a dataset with 1,117,217 training examples. We exclude HybridQA, InfoTabs, and Table Recognition from the training set to allow for held-out evaluation, leaving 1,031,082 examples in Tablet-M and 3,419,176 in the full dataset of our experiments. The capped distribution is shown in pie chart (b), Table 4 and further details on training sets are in Appendix 7.1.
@@
The model trained on Tablet-M performs competitively and in some cases, even better than Tablet-L (see Table 4, rows in gray), offering a compelling trade-off between dataset size and model effectiveness. That said, the full dataset still yields slightly better performance overall and remains a valuable resource for future work that can benefit from large-scale, lossless training.
The Benefit of Training on Table Interpretation Tasks Table Interpretation tasks represent 54.2% of the training examples in Tablet-L, making them the largest category by volume and the second longest in instruction length, following HybridQA (see Appendix 8.1). While deng2020turl demonstrated the benefits of these tasks for textual TU, it is unclear whether this carries over to VTU. If these tasks prove unhelpful, removing them would save significant resources.
@@
We therefore remove all Table Interpretation tasks from Tablet-M and fine-tune on this smaller version (we use Tablet-S as a shorthand for Tablet-Small) which excludes any tasks related to Column Type Annotation, Entity Linking, and Relation Extraction resulting in a dataset with a total of 690,467 training examples (67% of the example count in Tablet-M). As shown in Table 4, there are benefits to be gained from including these tasks. Models trained on Tablet-M outperform those trained on Tablet-S on six out of 8 held-in benchmarks (Table 3) and achieve comparable performance on the remaining two. Similar gains are observed in four out of five held-out tasks (Table 4). These findings demonstrate that the benefits observed in the textual domain (deng2020turl) extend to multimodal VTU.

### v3

@@ 3.4 Linking to Sources and Highlighting
Tablet is designed with extensibility in mind. Existing datasets such as TableInstruct (Zhang et al., 2024), DocStruct4M (Hu et al., 2024), and MMTab (Zheng et al., 2024b) do not provide a clear reference to the original examples from which their instructions were derived. However, the absence of such pointers, makes reuse, modification, and augmentation difficult. Additionally, image-based datasets such as MMTab lack serialized table versions in text format, providing only rendered images as representations. To address these issues, each example in Tablet includes both the identifier of the original example and the identifier of the corresponding table from the source dataset. Identifiers follow a simple, human-readable format, and the released code provides a function to retrieve the original examples given their IDs. To support future extensions, TABLET also provides tables serialized in HTML. For tables retrieved from Wikipedia, this corresponds to the raw table object from the article’s original HTML, which can be rendered using Wikipedia’s official CSS stylesheets or the rendering functions included in our code. All remaining tables are either provided as HTML in the seed datasets or converted into HTML from their serialized forms.
This feature is particularly useful in tasks involving cell pointing or highlighting. Prior work has shown that models can achieve competitive results by relying uniquely on highlighted values (An et al., 2022; Alonso et al., 2024). We also observed in our experiments that as long as highlighted cell values are mentioned in the prompt, models are able to perform the task (e.g., Column Type Annotation, Entity Linking, Relation Extraction, FeTaQA, and ToTTo) regardless of the table provided. We mitigate this, by exploiting the HTML serialization to directly locate the highlighted cells in the source tables. By cross-referencing the highlighted spans from the seed datasets with the actual table content, we generate versions of the tables with explicit highlights while preserving the original visual features.
@@ 3.5 Comparison with MMTab
MMTab (Zheng et al., 2024b) is also a dataset for visual table understanding; however, unlike Tablet, it does not include the full set of training examples from the original datasets (e.g., only 12.4% of ToTTo compared to our 91.8%; see Appendix 7.6), and it provides no development splits. MMTab covers 20 tasks while Tablet covers 22 (inculding instruction tuning); four tasks are unique to the former and five to the latter, and overall our dataset is 9.38 times larger (comprising 4,066,945 examples compared to MMTab’s approximate 433,376 examples). Moreover, only a few tasks in MMTab include identifiers that can be linked back to their source examples, limiting developers to the provided instruction text. While MMTab relies on synthetically rendered tables with predefined styles, which omit original formatting and embedded images, Tablet retains the original visualizations, enabling models to exploit a fuller range of visual features.
@@ 4 Experimental Setting
Our experiments are motivated by three questions: (1) Does training with original visualizations lead to more robust generalization to real-world tables, even on benchmarks that do not explicitly require visual features, or does the added visual information act as noise and degrade performance?; (2) Does fine-tuning on Tablet outperform existing resources on both seen and unseen VTU tasks?; and (3) Does fine-tuning on Tablet enable models to better integrate visual perception and table understanding for unseen Visual Table Question Answering tasks? We also examine whether training only with original visualizations performs better than mixing original and synthetic ones, whether a balanced task distribution is preferable to the full dataset distribution, and whether including various table interpretation tasks contributes to VTU performance.
Backbone Model All main experiments employ the same backbone model, namely Qwen2.5-VL with 7B parameters (Qwen et al., 2025). This model provides a good tradeoff between performance and computational requirements. Together with InternVL3 (Zhu et al., 2025), it was the best-performing open-weight VLM in its class, setting a new benchmark for Single Page Document VQA (Mathew et al., 2021) at the time of writing. Exploring the best backbone model for VTU or advanced SFT strategies is outside the scope of this work. Full SFT configurations are detailed in Appendix 8; results for other open-weight models on Tablet (test set) are reported in Appendix 8.3.
Evaluation We evaluate fine-tuned models on eight held-in tasks: three focus on Table QA (WikiTableQuestions, HiTabQA, FeTaQA), two on text generation (WikiBio, ToTTo), two on Table Numerical Reasoning tasks (TabMWP, TAT-QA), and one on Table NLI (TabFact). To test generalization, we further evaluate on seven held-out tasks: VisualTableQA, the new benchmark introduced in this work, HybridQA, InfoTabs, TabMCQ (Jauhar et al., 2016), AIT-QA (Katsis et al., 2022), Table Recognition, and PubHealthTab (Akhtar et al., 2022). The last four are directly extracted from MMTab and use instructions generated with their own templates, adding diversity in both table and instruction styles. Although included in Tablet, the 86,135 training examples from HybridQA, InfoTabs, and Table Recognition were excluded from fine-tuning to enable held-out evaluation.
WikiBio ToTTo WikiTQ HiTab TabFact FeTaQA Infotabs DegScore
Model org $\Delta$ org $\Delta$ org $\Delta$ org $\Delta$ org $\Delta$ org $\Delta$ org $\Delta$ total
0-Shot 2.5 +0.5 8.9 -1.9 50.8 -6.9 2.5 -8.6 59.4 -11.0 9.8 -0.1 64.5 -4.6 -28.90
@@
Table 2: Comparison of Qwen2.5-VL-7B-Instruct models fine-tuned on different Tablet-B variants and in zero-shot mode. We report performance when evaluating on original (lossless) table visualizations across held-in and held-out benchmarks and the corresponding degradation ($\Delta$) relative to evaluation on synthetic (lossy) tables. Negative $\Delta$ values indicate worse performance on original visualizations. We report exact match accuracy for WikiTQ, TabMWP, TabFact, and TAT-QA; BLEU for WikiBio, ToTTo, and FeTaQA; and F1 for HiTab (ToTTo on the dev set). Best results per setting (synth/org) are shown in bold. DegScore denotes the normalized aggregate degradation, computed separately for BLEU, accuracy, and F1 (see Appendix 8.2).
5 Results and Analysis
@@
[…]
Performance on VTU Tasks with Tablet We next evaluate whether Tablet improves performance on seen and unseen VTU tasks. While fine-tuned models are expected to perform better on tasks seen during training, improved performance on held-out datasets would suggest that training on Tablet enhances the model’s VTU capabilities. Tables 3 and 4 report our results on held-in and held-out tasks, respectively. We present comparisons between zero- and one-shot Qwen2.5-VL-7B [^4] We should note that Qwen2.5-VL-7B is a competitive baseline that has undergone extensive pre-training and instruction-tuning, achieving top scores in many TU and Visual Document Understanding benchmarks. and its fine-tuned instantiations trained on MMTab and different sizes of Tablet. For MMTab, all Wikipedia-sourced images are synthetic (61.8% of all tables in their dataset), whereas all Tablet versions use a mix of original (88%) and synthetic (12%) (See Appendix 7.2 for the distribution of image sources per task). For now, we focus solely on Tablet-L (a shorthand for Tablet-Large), which is our biggest dataset comprising 4M examples (we discuss smaller sizes in the next sections).
@@
Gemma 3 0-Shot 4.1 4.9 35.1 65.2 20.6 62.1 20.5 7.9
MMTab 4.6 18.1 29.7  73.2 7.6 56.6 21.3 6.9
[…]
Table 3: Evaluation on eight held-in datasets  with Qwen2.5-VL-7B-Instruct in 0/1-shot mode and fine-tuned on MMTab and different Tablet sizes. Results for Gemma 3 4B in 0-shot mode and fine-tuned on MMTab ,Tablet-M, and Tablet-L using LoRA are also included.Results for ToTTo correspond to evaluation on the development set. We report exact match accuracy for WikiTQ, TabMWP, TabFact, and TAT-QA; BLEU for WikiBio, ToTTo, and FeTaQA; and F1 for HiTab. Best performing model per task is shown in bold; we mark with $\ast$ models significantly different ($p<0.05$, using bootstrap resampling) from those trained on Tablet (highlighted in gray). Models that perform significantly better within the Tablet group are underlined. Results for other open-weight models on Tablet (test set) are reported in Appendix 8.3.
@@
Infotabs TabMCQ AIT-QA PubHealthTab HybridQA TabRec VTQA
0-Shot 64.5 84.4  51.7  64.7 34.2  24.5  42.4
MMTab  56.9 89.1  56.6  63.9 27.1  43.6  41.1
@@
Table 4: Left: Evaluation on five held-out datasets with Qwen2.5-VL-7B-Instruct (0-shot) and fine-tuned on MMTab and different Tablet sizes. We report Tree-Edit- Distance-based Similarity for Table Recognition (TabRec) and accuracy for all other benchmarks. HybridQA results are reported on the development set. Best performing model per task is shown in bold; we mark with $\ast$ models significantly different ($p<0.05$, using bootstrap resampling) from those trained on Tablet (highlighted in gray). Models that perform significantly better within the Tablet group are underlined. Right: Distribution of examples per task in Tablet-L (b) and Tablet-M (a).
@@
We observe that the model fine-tuned on Tablet-L  performs significantly better on held-in and held-out datasets. It dominates in all QA benchmarks except HybridQA, where long-context proficiency is likely reduced since its training set was excluded from fine-tuning. Notably, in tasks such as HiTab, FeTaQA, and TAT-QA, where Tablet includes fewer training examples than MMTab, the model still performs considerably better, suggesting benefits from transfer across tasks, original visualizations, or simply due to a higher-quality subset of examples. Finally, tasks that include hierarchical table structures such as ToTTo and HiTab show clear gains. Table cell hierarchy is not lost in synthetic tables, but is often conveyed more clearly in their original visualizations. Our results are encouraging, particularly when considering that portions of the held-out datasets, including test sets, may have been encountered during pretraning. Models trained on Tablet outperform 0/1-shot Qwen2.5-VL-7B on 12 out 14 (held-in and held-out) tasks. Interestingly, our gains extend beyond unseen tasks. For instance, ToTTo, HiTab, TabFact, FeTaQA, and TAT-QA contain the same number of training examples across all Tablet variants, yet performance consistently improves as additional tasks are incorporated. This suggests that fine-tuning on a broader task set facilitates transfer learning, yielding improvements even on tasks where the model was already competitive.
@@
Optimal Tablet Size Since training on 4M examples is resource-intensive, and tasks in Tablet-L exhibit substantial variation in size (see (b) pie chart in Table 4), we evaluate whether a smaller but more balanced dataset might perform comparably. We create Tablet-M (a shorthand for Tablet-Medium) by capping each task at 140k examples. For benchmarks exceeding this cap (Column Type Annotation, Entity Linking, WikiBio), a random 140k subset is sampled and included as representative for that task. This results in a dataset with 1,117,217 training examples. We exclude HybridQA, InfoTabs, and Table Recognition from the training set to allow for held-out evaluation, leaving 1,031,082 examples in Tablet-M and 3,419,176 in the full dataset of our experiments. The capped distribution is shown in pie chart (b) and Table 4 (further details on training sets are in Appendix 7.1).
@@
Models trained on Tablet-M perform competitively and in some cases, even better than Tablet-L (see Table 4, rows in gray), offering a compelling trade-off between dataset size and model effectiveness. That said, the full dataset still yields slightly better overall performance when fully fine-tuning a larger model and remains a valuable resource for future work that can benefit from large-scale, lossless training.
Robustness Across Models and Training Regimes We observe similar performance trends when training Gemma 3 4B with LoRA (Table 3, Appendix 12). Fine-tuning on Tablet-L yields substantial gains over the zero-shot baseline on six of eight tasks, with the largest improvements on visually rich datasets like HiTab (+22.4). Notably, both Tablet-M and Tablet-L significantly outperform MMTab on WikiBio, ToTTo, HiTab, and FeTaQA, despite using parameter-efficient LoRA rather than full fine-tuning employed for Qwen2.5-VL. This demonstrates that Tablet’s benefits are robust to model architecture, model scale, and training methodology.
@@
The Benefit of Training on Table Interpretation Tasks Table Interpretation tasks represent 54.2% of the training examples in Tablet-L, making them the largest category by volume and the second longest in instruction length, following HybridQA (see Appendix 8.1). While Deng et al. (2020) demonstrated the benefits of these tasks for textual TU, it is unclear whether this carries over to VTU. If these tasks prove unhelpful, removing them would save significant resources. We therefore remove all Table Interpretation tasks from Tablet-M and fine-tune on this smaller version (we use Tablet-S as a shorthand for Tablet-Small) which excludes any tasks related to Column Type Annotation, Entity Linking, and Relation Extraction resulting in a dataset with a total of 690,467 training examples (67% of the example count in Tablet-M). As shown in Table 4, there are benefits to be gained from including these tasks. Models trained on Tablet-M outperform those trained on Tablet-S on six out of 8 held-in benchmarks (Table 3) and achieve comparable performance on the remaining two. Similar gains are observed in four out of five held-out tasks (Table 4). These findings demonstrate that the benefits observed in the textual domain (Deng et al., 2020) extend to multimodal VTU.
Performance on Unseen Visual Table Question Answering Table 4 also reports results on VisualTableQA (VTQA), our benchmark designed to evaluate models’ ability to jointly reason over visual cues and tabular structure. Fine-tuning on Tablet yields substantial improvements over both zero-shot (42.4%) and MMTab (41.1%) baselines, with Tablet-M achieving the best performance (47.8%). This represents a 13% relative improvement over zero-shot and 16% over MMTab, demonstrating that exposure to Tablet’s visually rich original table images enables models to develop compositional skills that transfer to this challenging unseen task. Notably, these gains emerge without any task-specific VisualTableQA examples in the training data, confirming that Tablet’s diverse task mixture and authentic visualizations foster generalizable visual reasoning capabilities. Results are averaged over three random seeds and reported as exact match accuracy.
Visual Complexity Analysis We hypothesize that Tablet is especially effective for visually complex tables; those deviating from regular, monochromatic layouts toward richer structures with stylistic cues, embedded images, irregular formatting, and diverse styling. These tables suffer most from serialization and motivate our work. Since Tablet contains both format-aware serialized representations (e.g., HTML) and rendered visual forms, we compute a visual-complexity score aggregating factors such as row/column-span irregularity, visual entropy, font diversity, embedded image features, and edge irregularity (see Appendix 10 for details). We rank tables by this score and group them into 20 equally spaced bins for analysis, omitting bins with fewer than 5 examples for clarity. Figure 2 shows results for tasks with the most pronounced model differences (see Appendix 11 for all benchmarks). Across these tasks, models trained on Tablet remain robust as visual complexity increases. For tasks like TABMWP, fine-tuning on table images proves crucial to avoid performance degradation, with Tablet-trained models showing consistently higher performance overall.
0-Shot MMTab   Tablet-S Tablet-M Tablet-L Refer to caption
Figure 2: Model (Qwen2.5-VL-7B-Instruct) performance across different levels of table visual complexity (x axis). Higher levels correspond to more visually complex tables. Complexity levels are comparable across tasks, e.g., level 4 represents the same complexity range in every benchmark.